# Nonlinear Bayesian Filtering with Natural Gradient Gaussian Approximation

### v1

@@ VIII-B Nonlinear System Simulation: Air-traffic Control
Next, we perform simulations for a nonlinear system, focusing on an air-traffic control scenario where an aircraft executes a maneuvering turn in a horizontal plane at a height $h=50$ with respect to the radar [14]. The state vector is defined as $x=\left[p_{x},\ v_{x},\ p_{y},\ v_{y},\ \omega\right]^{\top}$, where $p_{x},p_{y}$ are the positions, $v_{x},v_{y}$ are the velocities in the planar coordinates, and $\omega$ is the turn rate. The kinematics of the turning motion, represented by the state transition model, and the measurement model are given by:
$x_{t+1}$ $=\begin{bmatrix}1&\frac{\sin\omega\Delta t}{\omega}&0&-\frac{(1-\cos\omega\Delta t)}{\Omega}&0\\ 0&\cos\omega\Delta t&0&-\sin\omega\Delta t&0\\ 0&\frac{(1-\cos\omega\Delta t)}{\omega}&1&\frac{\sin\omega\Delta t}{\omega}&0\\ 0&\sin\omega\Delta t&0&\cos\omega\Delta t&0\\ 0&0&0&0&1\end{bmatrix}x_{t}+\xi_{t},$
$y_{t}$ $=\begin{bmatrix}\sqrt{p_{x}^{2}+p_{y}^{2}+h^{2}}\\ \operatorname{atan}\left(\frac{p_{y}}{p_{x}}\right)\\ \operatorname{atan}\left(\frac{h}{\sqrt{p_{x}^{2}+p_{y}^{2}}}\right)\\ \frac{p_{x}\dot{p}_{x}+p_{y}\dot{p}_{y}}{\sqrt{p_{x}^{2}+p_{y}^{2}+h^{2}}}\end{bmatrix}+\zeta_{t}.$
@@
Figure 4: Box plot of RMSE for EKF, UKF, IEKF, PLF, NANO filter, and its robust variants: $\beta$-NANO filter, Huber-NANO filter, and Weight-NANO filter with various parameter values, for the air-traffic model with measurement outliers.
Similar to the linear case, we validate our method under two scenarios for the nonlinear system. In the first scenario, shown in Fig. 3, the NANO filter is compared with the EKF, UKF, IEKF, and PLF. The results demonstrate that the NANO filter significantly outperforms these established filters, achieving nearly a half reduction in RMSE. This substantial improvement highlights the NANO filter’s superior effectiveness in handling nonlinear dynamics. In the second scenario, we introduce measurement outliers to reflect conditions where the measurement data deviates from the assumed model. Specifically, we assume the actual measurement noise follows a mixture model: $\zeta_{t}\sim 0.9\cdot\mathcal{N}(\zeta_{t};0,R)+0.1\cdot\mathcal{N}(\zeta_{t};0,100\cdot R)$, indicating a 10% probability that measurements are contaminated by noise with a tenfold increase in standard deviation. As shown in Fig. 4, robust variants of the NANO filter—including the $\beta$-NANO, Huber-NANO, and Weight-NANO filters—significantly improve estimation accuracy in the presence of measurement outliers. Notably, the $\beta$-NANO filter with $\beta=10^{-2}$ exhibits the best performance, effectively mitigating the impact of outliers.
@@ VIII-C Real-world Experiment: Unmanned Ground Vehicle Localization
Figure 5: (a) The UGV. (b) Experiment field. The three red traffic cones serve as landmarks for positioning.
@@
where $\Delta t=0.0667s$ is the sample period, $v_{t}$ is the longitudinal velocity, $\omega_{t}$ is the yaw rate, and $\xi_{t}$ is the process noise. The observation model is given by:
$y_{t}=[d^{1}_{t},d^{2}_{t},d^{3}_{t},\alpha^{1}_{t},\alpha^{2}_{t},\alpha^{3}_{t}]^{\top}+\zeta_{t},$
@@
$d^{i}$ $=\|(p^{tc,i}_{x}-p_{x}-l\cos\theta,p^{tc,i}_{y}-p_{y}-l\sin\theta)\|_{2},$ (51)
$\alpha^{i}$ $=\arctan\left(\frac{p^{tc,i}_{y}-p_{y}-l\sin\theta}{p^{tc,i}_{x}-p_{x}-l\cos\theta}\right)-\theta.$
Here $(p^{tc,i}_{x},p^{tc,i}_{y})$ is the position of the $i$-th traffic cone, and $l$ represents the longitudinal installation offset of the Lidar with respect to the robot center, as depicted in Fig. 6.
@@
Figure 8: A sample of the real trajectory and its estimates of the UGV estimation.
The experiment began by collecting raw data, including ground truth states, control inputs, and Lidar point clouds. The Lidar data were processed to extract relative distances and angles to the traffic cones, resulting in a dataset comprising trajectories totaling 11 minutes. Using ground truth data obtained from a high-precision motion capture system and the measurement data, we estimated noise parameters, assuming Gaussian distributions with independent noise components. The dataset was then segmented into 50 sub-trajectories, each lasting 6.67 seconds (100 time steps). Figures 7 and 8 compare the RMSE performance of the EKF, UKF, IEKF, PLF, and NANO filters, including their robust variants. The results indicate that the NANO filter consistently outperforms traditional Gaussian filters, achieving higher accuracy. Furthermore, the robust NANO variants demonstrate improved performance, highlighting their effectiveness in handling measurement outliers typically encountered in real-world conditions.
@@
Previous results for simulations and experiment demonstrate that the NANO filter significantly improves RMSE performance compared to traditional Gaussian filters. However, a critical aspect to consider is the computational cost per time step for each algorithm. Table II shows that for nonlinear systems, the average runtime of NANO filter is approximately twice that of the EKF, which indicates that despite being an iterative method, NANO maintains high efficiency and quickly converges. This balance of improved accuracy with reasonable computational demands highlights the practicality of the NANO filter, even in real-time applications.
@@ IX Conclusion
In this paper, we address the estimation errors commonly introduced by linearization techniques in Gaussian filters like the EKF and UKF for nonlinear systems. We reformulate the prediction and update steps of Gaussian filtering as optimization problems. While the prediction step mirrors moment-matching filters by calculating the first two moments of the prior distribution, the update step poses more challenges due to its highly nonlinear nature. To overcome these issues, we propose an iterative approach called the NANO filter, which avoids linearization by directly optimizing the update step using the natural gradient derived from the Fisher information matrix. This allows us to account for the curvature of the parameter space and ensures more accurate updates. We prove that the NANO filter converges locally to the optimal Gaussian approximation at each time step and show that it provides exponential error bounds for nearly linear systems with low noise. Experimental results demonstrate that the NANO filter outperforms popular filters like EKF, UKF, and others, while maintaining similar computational efficiency.
@@
This proof is inspired by Section 6.2.2 in [45]. According to (16) and (26), we have
@@
According to (25) and (26), we have
$\left.\frac{\partial J}{\partial\hat{x}_{t}}\right|_{v^{(i)}}$ $=-\left(P_{t}^{-1}\right)^{(i+1)}\delta\hat{x}_{t}^{(i)},$ (53)
$\left.\frac{\partial J}{\partial P_{t}^{-1}}\right|_{v^{(i)}}$ $=-\frac{1}{2}\left(P_{t}\right)^{(i)}\delta P_{t}^{(i)}\left(P_{t}\right)^{(i)}.$

### v2

@@ VIII-B Nonlinear System Simulation: Air-traffic Control
Next, we perform simulations for a nonlinear system, focusing on an air-traffic control scenario where an aircraft executes a maneuvering turn in a horizontal plane at a height $h=50$ with respect to the radar [14]. The state vector is defined as $x=\left[p_{x}\ v_{x}\ p_{y}\ v_{y}\ \omega\right]^{\top}$, where $p_{x},p_{y}$ are the positions, $v_{x},v_{y}$ are the velocities in the planar coordinates, and $\omega$ is the turn rate. The state transition model representing the kinematics of the turning motion and the measurement model are given by:
$x_{t+1}$ $=\begin{bmatrix}1&\frac{\sin\omega_{t}\Delta t}{\omega_{t}}&0&-\frac{1-\cos\omega_{t}\Delta t}{\Omega_{t}}&0\\ 0&\cos\omega_{t}\Delta t&0&-\sin\omega_{t}\Delta t&0\\ 0&\frac{1-\cos\omega_{t}\Delta t}{\omega_{t}}&1&\frac{\sin\omega_{t}\Delta t}{\omega_{t}}&0\\ 0&\sin\omega_{t}\Delta t&0&\cos\omega_{t}\Delta t&0\\ 0&0&0&0&1\end{bmatrix}x_{t}+\xi_{t},$
$y_{t}$ $=\begin{bmatrix}\sqrt{p_{x,t}^{2}+p_{y,t}^{2}+h^{2}}\\ \operatorname{atan}\left(\frac{p_{y,t}}{p_{x,t}}\right)\\ \operatorname{atan}\left(\frac{h}{\sqrt{p_{x,t}^{2}+p_{y,t}^{2}}}\right)\\ \frac{p_{x,t}\dot{p}_{x,t}+p_{y,t}\dot{p}_{y,t}}{\sqrt{p_{x,t}^{2}+p_{y,t}^{2}+h^{2}}}\end{bmatrix}+\zeta_{t}.$
@@
Figure 4: Box plot of RMSE for EKF, UKF, IEKF, PLF, NANO filter, and its robust variants with various parameter values, for the air-traffic model with measurement outliers.
Here, $q_{1}=2.5$ and $q_{2}=10^{-6}$. Similar to the linear case, we validate our method under two scenarios for the nonlinear system. In the first scenario, shown in Fig. 3, the NANO filter is compared with EKF, UKF, IEKF, and PLF. The results demonstrate that NANO filter significantly outperforms these established filters, achieving nearly a half reduction in RMSE. This substantial improvement highlights the NANO filter’s superior effectiveness in handling nonlinear dynamics. In the second scenario, we introduce measurement outliers to reflect conditions where the measurement data deviates from the assumed model. Specifically, we assume the actual measurement noise follows a mixture model: $\zeta_{t}\sim 0.9\cdot\mathcal{N}(\zeta_{t};0,R)+0.1\cdot\mathcal{N}(\zeta_{t};0,100\cdot R)$, indicating a 10% probability that measurements are contaminated by noise with a tenfold increase in standard deviation. As shown in Fig. 4, robust variants of the NANO filter—including the $\beta$-NANO, Huber-NANO, and Weight-NANO filters—significantly improve estimation accuracy in the presence of measurement outliers. Notably, the $\beta$-NANO filter with $\beta=10^{-2}$ exhibits the best performance, effectively mitigating the impact of outliers.
@@ VIII-C Real-world Experiment: Unmanned Ground Vehicle Localization
Figure 5: The UGV and the experiment field. The three red traffic cones serve as landmarks for positioning.
@@
where $\Delta t=0.0667$ is the sample period, $v_{t}$ is the longitudinal velocity, $\omega_{t}$ is the yaw rate, and $\xi_{t}$ is the process noise. The measurement model is given by:
$y_{t}=[d^{1}_{t}\ d^{2}_{t}\ d^{3}_{t}\ \alpha^{1}_{t}\ \alpha^{2}_{t}\ \alpha^{3}_{t}]^{\top}+\zeta_{t},$
@@
$d^{i}$ $=\|(p^{\mathrm{tc},i}_{x}-p_{x}-l\cos\theta,p^{\mathrm{tc},i}_{y}-p_{y}-l\sin\theta)\|_{2},$ (51)
$\alpha^{i}$ $=\arctan\left(\frac{p^{\mathrm{tc},i}_{y}-p_{y}-l\sin\theta}{p^{\mathrm{tc},i}_{x}-p_{x}-l\cos\theta}\right)-\theta.$
Here $(p^{\mathrm{tc},i}_{x},p^{\mathrm{tc},i}_{y})$ is the position of the $i$-th traffic cone, and $l$ represents the longitudinal installation offset of the Lidar with respect to the robot center, as depicted in Fig. 6.
@@
Figure 8: Two samples of the real trajectories and their estimates.
The experiment begin by collecting raw data, including ground truth states from the high-precision motion capture system, control inputs, and Lidar point clouds. The Lidar data are processed to extract relative distances and angles to the traffic cones, resulting in a dataset comprising trajectories totaling 11 minutes. The noises are assumed to be Gaussian, whose parameters are estimated from the ground truth and measurement data using maximum likelihood estimation. The dataset is then segmented into 50 sub-trajectories, each lasting 6.67 seconds (100 time steps). Figures 7 and 8 compare the RMSE performance of the EKF, UKF, IEKF, PLF, and NANO filters, including their robust variants. The results indicate that the NANO filter consistently outperforms traditional Gaussian filters, achieving higher accuracy. Furthermore, the robust NANO variants demonstrate improved performance, demonstrating their effectiveness in handling measurement outliers typically encountered in real-world conditions.
@@
Previous results for simulations and experiment demonstrate that the NANO filter significantly improves RMSE performance compared to traditional Gaussian filters. However, a critical aspect to consider is the computational cost per time step for each algorithm. Table II shows that for nonlinear systems, the average runtime of NANO filter is approximately 2 to 3 times that of the EKF, which indicates that despite being an iterative method, NANO filter maintains high efficiency and quickly converges. This balance of improved accuracy with reasonable computational demands highlights the practicality of the NANO filter, even in real-time applications.
@@ IX Conclusion
In this paper, we address the estimation errors commonly introduced by linearization techniques in Gaussian filters like the EKF and UKF for nonlinear systems. We reformulate the prediction and update steps of Gaussian filtering as optimization problems. While the prediction step mirrors moment-matching filters by calculating the first two moments of the prior distribution, the update step poses more challenges due to its highly nonlinear nature. To overcome these issues, we propose an iterative approach called the NANO filter, which avoids linearization by directly optimizing the update step using the natural gradient derived from the Fisher information matrix. This allows us to account for the curvature of the parameter space and ensures more accurate updates. We prove that the NANO filter converges locally to the optimal Gaussian approximation at each time step and show that it provides exponential error bounds for nearly linear measurement models with low noise. Experimental results demonstrate that the NANO filter outperforms popular filters like EKF, UKF, and others, while maintaining comparable computational efficiency.
@@
This proof is inspired by Section 6.2.2 in [44]. According to (16) and (26), the second-order derivative of $J(\hat{x}_{t},P_{t})$ with respect to $\hat{x}_{t}$ satisfies
@@
According to (25), we have
$\left.\frac{\partial J}{\partial\hat{x}_{t}}\right|_{v^{(i)}}$ $=-\left(P_{t}^{-1}\right)^{(i+1)}\delta\hat{x}_{t},$ (53)
$\left.\frac{\partial J}{\partial P_{t}^{-1}}\right|_{v^{(i)}}$ $=-\frac{1}{2}\left(P_{t}\right)^{(i)}\delta P_{t}^{-1}\left(P_{t}\right)^{(i)}.$